Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 178     Entry time: 05 May 2023, 06:44
Author: Tokushi Sato 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 12.4 keV 200 pulses / train, relatively stable intensity ~0.8 mJ / pulse

Optical laser delivery

  • 420 nm PP laser
  • ns laser for jet imaging

Achievements/Observations

  • M1 & M2 offset mirror optimized again (not really optimum but acceptable)
  • NKB focus checked and defocused to have about 2um beam
  • Collected a few dark runs and about 30 runs with PP laser in adaptive gain mode

Issues

  • AGIPD Fixed MG mode issue
  • PP-laser sync issue, but it was solved quickly, thanks to Lbsync team
  • Beam pointing in the FEL Imager was quite off in the beginning of the shift 
  • Hit rate too low (~0.3%)
ELOG V3.1.4-7c3fd00